OpenStudy (lgbasallote):

Three people are issued movie tickets with seat assignments, however, it was dark and the three people decide to just sit on the three seats without ascertaining if it is the correct numbered seat. What is the probability they are all sitting on the correct seats?

OpenStudy (anonymous):

I would assume we would need to know how many seats are there O_o .

OpenStudy (anonymous):

I would also assume that you would need to know then umberof seats

OpenStudy (anonymous):

A lot needs to be thought off. You need to take into account the probability of the first guy of choosing the right seat in order for the second. I would say 1/6 is the answer.

OpenStudy (anonymous):

1/3 chance of the first times 1/2 chance for the second?

OpenStudy (anonymous):

@ChmE . How? You know nothing except that 3 people bough tickets. The probability changes as the number of seats changes.

OpenStudy (anonymous):

Isn't it: 1/3!

OpenStudy (anonymous):

I assumed the number of seats is 3

OpenStudy (anonymous):

decide to just sit on the three seats. When I read this I interpret it as they see the 3 seets they bought but don't know their individual seat assignment

OpenStudy (anonymous):

@sauravshakya Why? You don't know how many seats are there.

OpenStudy (anonymous):

I guess. I am not sure if you can make that assumption though...

OpenStudy (anonymous):

So the first guy walks up and chooses. He has a 1/3 chance. The second walks up and he now has a 1/2 chance. The last is grandfathered in given the first two were crrect

OpenStudy (anonymous):

Because they have three tickets of three sits and the question said that they sit on the three sits.

OpenStudy (anonymous):

What if there was 50 seats? it would be a 1/50 chance.

OpenStudy (anonymous):

3/50 sorry.

OpenStudy (anonymous):

1/50 time 1/49 times 1/48

OpenStudy (anonymous):

If there was 50 sits then 1/50P3

OpenStudy (anonymous):

Same as @ChmE did

OpenStudy (anonymous):

I guess I just have to be intuitive in this question i guess :/ .

OpenStudy (anonymous):

My interpretation of the problem is they bought 3 tickets. They can't see the individual seat assignment illustrated by "without ascertaining if it is the correct numbered seat" But they do see THE 3 seats ilustrated by "three people decide to just sit on THE three seats"

OpenStudy (anonymous):

Ohh. If it just said "on" instead of "the" the question would change completely.

OpenStudy (anonymous):

I get it.

OpenStudy (anonymous):

@lgbasallote you've been silent. What are your thoughts?

OpenStudy (lgbasallote):

sorry i was away

OpenStudy (lgbasallote):

anyway..the number of seats doesn't matter

OpenStudy (anonymous):

Wouldnt the probably just be then 1/nP3 ?

OpenStudy (anonymous):

Do you know the answer or are you looking for help with this solution

OpenStudy (anonymous):

Where n is the number of seats?

OpenStudy (lgbasallote):

yes i do @ChmE

OpenStudy (anonymous):

My answer is 1/6. Correct or no?

OpenStudy (lgbasallote):

why so?

OpenStudy (anonymous):

I've already explained it above

OpenStudy (lgbasallote):

you were inferencing above that the number of seats affect the outcome

OpenStudy (anonymous):

It has to.

OpenStudy (shane_b):

@ChmE's answer should be correct... 1 in 6. The question doesn't provide the number of seats so it's assumed to be 3.\[\frac{1}{3}*\frac{1}{2}=\frac{1}{6}\]

OpenStudy (anonymous):

Chance of first guy chooses correct 1/3. Given he is correct, there are 2 seats left and the chance for the second guy to be correct is 1/2. Given he is also correct, the last guy has a 1/1 chance of being correct

OpenStudy (shane_b):

If the number of seats were higher, it would obviously affect the odds of each selection above.

OpenStudy (lgbasallote):

oh so that's why

OpenStudy (anonymous):

I got a question for you now @lgbasallote

OpenStudy (lgbasallote):

shoot

OpenStudy (shane_b):

@lgbasallote: What if the number of available seats was 5?\[\frac{1}{5}*\frac{1}{4}*\frac{1}{3}=\frac{1}{60}\]

OpenStudy (shane_b):

Just to illustrate...

OpenStudy (lgbasallote):

why not include 1/2?

OpenStudy (anonymous):

You are on a game show. There are 3 doors. Behind one of the doors is a car. You choose door number 1. The host walks up to door 2 and reveals nothing behind it. He asks you if you want to change your guess. Is it in your best interest to change doors?

OpenStudy (lgbasallote):

that's monty hall's problem

OpenStudy (anonymous):

Ohh I remember that one. Behind two doors is a goat. Behind 1 is a car.

OpenStudy (lgbasallote):

that urban myth states that you switch

OpenStudy (anonymous):

there is no myth. Just an extra 33% chance of getting it correct if you switch.

OpenStudy (anonymous):

It's shown in the movie "21"

OpenStudy (lgbasallote):

all math for me is urban myth

OpenStudy (anonymous):

haha

OpenStudy (shane_b):

At that point in time you have a 50/50 chance of choosing the right door...switching doors won't change your odds *at that point in time*.

OpenStudy (anonymous):

Wrong. In the beginning you have 1/3 chance is in door 1 and 2/3 chance of being in door 2 and 3. After he reveals door 2 is wrong, all 2/3 chance gets put on door 3. It is accounting for variable change

OpenStudy (lgbasallote):

u.r.b.a.n m.y.t.h

OpenStudy (anonymous):

I guess if you say at that exact point in time looking at only those 2 doors and forgetting there were 3. You would be correct

OpenStudy (shane_b):

You won't be able to convince that switching doors will increase your chances at that point in time. It depends on how you look at it...but ultimately it boils down to a 50/50 chance that you were right at that point.

OpenStudy (anonymous):

It's a tough concept to grasp, but in my stats class a couple years ago, we spent a couple days talking about these types of problems.

OpenStudy (shane_b):

Yea, I've dealt with it before as well. It's like trying to roll, say 23, on a theoretical die with 100 sides. With each successive roll your overall odds of rolling a 23 goes up in theory. However, each roll is a completely separate event with the same exact odds of hitting a 23.

OpenStudy (shane_b):

The (false) assumption here is that previous rolls affect future rolls.

OpenStudy (anonymous):

The die example you gave is independent events. My example the events are not independent

OpenStudy (shane_b):

How are they not independent in your example?

OpenStudy (shane_b):

I see it as going from 1 in 3 to 1 in 2...so I guess you could say they are not independent.

OpenStudy (anonymous):

You chance of door 1 is 1/3. He tells you door 2 is wrong and you see, oh hey, if door 2 is wrong and my chance of 2 or 3 was 2/3. Then because of the event where he said door 2 is wrong, door 3 now has all the 2/3 chance. i should switch.